OpenStudy (anonymous):

factor completely, if polynomial is not factorable write prime A. b^2 -16b +16 B. 10c^2 -17c +3 C. p^2+2p-24 D. c^3 + 8

OpenStudy (mathmale):

Hi. haven't you already posted the first of these four problems? I gave you feedback on A.

OpenStudy (anonymous):

I know but my computer started acting up

OpenStudy (anonymous):

could you help

OpenStudy (mathmale):

b^2 -16b +16 has the coefficients a=1, b=-16 and c=16. Look up the quadratic formula. Substitute these values of a, b and c into that formula. This will give you the roots of b^2 -16b +16.

OpenStudy (anonymous):

what quadratic formula

OpenStudy (mathmale):

Cookie: the quadratic formula is really important in algebra. If you haven't already seen it, you will see it very soon. If y ou have seen it before, please review it.

OpenStudy (anonymous):

i havent

OpenStudy (mathmale):

b^2 -16b +16 can be factored, but the factors are not 'pretty.' I've used the quadratic formula to find the two 'roots' or 'solutions' of b^2 -16b +16 = 0.

OpenStudy (mathmale):

Are you in an algebra class or are you studying math online?

OpenStudy (anonymous):

class

OpenStudy (mathmale):

I'd suggest you ask your teacher tomorrow whether or not you can use the quadratic formula to find the roots of b^2 -16b +16 = 0, and thus find the factors. b^2 -16b +16 = 0 is not easily factored without the q. formula.

OpenStudy (anonymous):

I dont think we can

OpenStudy (anonymous):

Could you help with the next one though

OpenStudy (mathmale):

The quadratic formula looks like\[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

b^2 -16b +16 has the coefficients a=1, b=-16 and c=16.

OpenStudy (mathmale):

Substitute these values into the quadratic formula: a=1, b=-16 and c=16.

OpenStudy (mathmale):

Solve for x.

OpenStudy (mathmale):

Example: Were you to obtain the two roots x and y, then the corresponding factors would be (b-x) and (b-y).

OpenStudy (anonymous):

could we skip a for now

OpenStudy (mathmale):

I'm sorry, Cookie, but I have to get off the computer. If you continue working on this problem, tag another moderator for help, or wait 'til I return later this evening.

OpenStudy (mathmale):

This would be a good time for us to stop. My suggestion is that you post (b) as a new question. Someone will surely respond.

OpenStudy (mathmale):

Forgive me, but I'm saying 'bye' for the present.

OpenStudy (anonymous):

k